Shift summary Day 23 November 1999.

Nickolas Solomey (solomey@hep.uchicago.edu)
Tue, 23 Nov 1999 15:29:01 -0600

Shift Summary for Day shift, 23 Nov 1999

On Shift: Nick

Hours of beam available at all intensities: 0.0
Hours of beam available at nominal intensity: 0.0
Hours of E799 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time: Scheduled down time for RunII studies
========================================================================
Runs ended this shift: none
========================================================================
Tapes brought to FCC this shift: none.
========================================================================
Accesses during shift: Yes many, see shift description below.
========================================================================
Overview of shift: Cosmic data taking, when we were not changing a csi
channel or other DA problem, Julie is working on it.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15034 6:55 Ongoing cosmic run from previous shift
9:30 Hogan and Nick made access to rewire HV for DC4.
9:45 Hogan increased DC1 by 30 volts.
10:00 Cleaned tape drives.
10:20 Tweaked E_{tot} trigger thresholds.
11:00 Doug and Jason #1 made access to do a z survey.
11:50 Julie and Jason #2 went on access for a couple of
pmt swaps, channels 1096 and 2893.
12:30 Leon Beverely went on access escorted by Nick and
Leo to put in a AC stability monitoring system.
13:20 Finished labeling a lot of tapes, should last
two weeks at least.
14:00 Julie comes out of access and says no triggers
there is a problem she has to return to fix
with Bob T at 2:30. It might be a while before
triggers are running again.
14:30 Valeri, Leo and Sasha go on access to work on VV'
14:50 Bob T. and Julie go on access to work on pipeline.
15:05 BOb T. and Julie out of access, triggers working.
15:15 CsI monitor is not updating but DC is, so as
soon as Valeri is done we need to end the run
and do a dart fresh start (Julie said to summaries
it as she is unhappy but Jason is happy!, I have
no comments to add).
15:30 end of shift.

=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=
Dr. Nickolas Solomey, Research Scientist, Enrico Fermi Institute
The University of Chicago, 5640 S. Ellis Ave, Chicago, Ill. 60637 USA
Phone: (773) 702-3278, fax: x-6317, http://hep.uchicago.edu/~solomey